Mr Young bought 2 dozen eggs. He used 1/2 of them to bake a cake, 1/3 of the remainder to bake some cookies and the rest for some tarts. Find the ratio of the number of eggs used for the cake to the number of eggs used for the cookies to the number of eggs used for the tarts.
Answer by ikleyn(52803)

.
Mr Young bought 2 dozen eggs. He used 1/2 of them to bake a cake,
1/3 of the remainder to bake some cookies and the rest for some tarts.
Find the ratio of the number of eggs used for the cake to the number of eggs
used for the cookies to the number of eggs used for the tarts.
~~~~~~~~~~~~~

2 dozen eggs = 24 eggs.


Cake = 1/2 of 24, i.e. 12 eggs.


The remainder = 24 - 12 = 12 eggs.


Some cookies = 1/3 of the remainer eggs = 1/3 of 12 eggs = 4 eggs.


Some tarts = 12 - 4 = 8 eggs.


The ratio cake : cookies : tarts = 12 : 4 : 8 = (canceling/reducing common factor 4) = 3 : 1 : 2.    ANSWER

Solved.